1. The pH of an aqueous solution at 25°C was found to be 4.60.  

The pOH of this solution is .  

The hydronium ion concentration is  M.

The hydroxide ion concentration is  M.

2.The pOH of an aqueous solution at 25°C was found to be 3.90.  

The pH of this solution is .  

The hydronium ion concentration is  M.

The hydroxide ion concentration is  M.

3.The hydronium ion concentration in an aqueous solution at 25°C is 2.2×10-2M.

The hydroxide ion concentration is  M.

The pH of this solution is .

The pOH is .

4.The hydroxide ion concentration in an aqueous solution at 25°C is 2.2×10-2 M.

The hydronium ion concentration is  M.  

The pH of this solution is .

The pOH is .
Answer #1

pOH=14-pH=14-4.60=9.4

pH=-logH+